### v2

# Symmetry-Preserving Finite-Difference Schemes and Auto-Bäcklund Transformations for the Schwarz Equation

Abstract
It is demonstrated that one of the equations from the Lie classification list of second-order ODEs is a first integral of the Schwarz equation. As symmetry-preserving finite-difference schemes have been previously constructed for both equations, the preservation of a similar connection between these schemes is studied. It is shown that the schemes for the Schwarz equation and the second-order ODE can be related through a Bäcklund-type difference transformation.
In addition, previously unexamined aspects of the difference scheme for the second-order ODE are discussed, including its singular solution and the complete set of difference first integrals.
keywords:
Lie point symmetry , numerical scheme , Schwarz equation , Bäcklund transformation , first integral
MSC:
[2020] 65L12 , 34C14
journal: Springer Proceedings
[…]
[label1]organization=School of Mathematics & Geoinformatics, Institute of Science, Suranaree University of Technology, city=Nakhon Ratchasima, postcode=30000, country=Thailand
[…]
[label2]organization=Keldysh Institute of Applied Mathematics, Russian Academy of Sciences, addressline=Miusskaya Pl. 4, city=Moscow, postcode=125047, country=Russia
1 Inroduction
Symmetries are closely related to the geometric properties of equations and play a crucial role in studying problems in mathematical physics [1, 2]. Knowledge of the symmetry of a differential equation often allows one to identify its conservation laws and exact solutions, as well as to reduce its order. As a particular case, for ordinary differential equations (ODEs), symmetries provide a means to obtain first integrals (using Noether’s theorem [3], the adjoint equation method, or the direct method [4]). In some cases, when a sufficient number of first integrals is known, an ODE can be completely integrated through algebraic calculations alone. Symmetry analysis also facilitates the reduction of higher-order equations to well-known reference equations of the same order or the lowering of their order by reduction on a subgroup. Among the numerous examples of such reductions, here we would like to highlight the works [5, 6] of Professor Kilkothur M. Tamizhmani and co-authors on the symmetries and reductions of third- and second-order ODEs to known cases, as well as the study of their Painlevé properties.
In this paper, we consider two particular ODEs and their discrete symmetry-preserving analogues. The first ODE is the Schwarz equation
$\frac{y^{\prime\prime\prime}}{y^{\prime}}-\frac{3}{2}\left(\frac{y^{\prime\prime}}{y^{\prime}}\right)^{2}=0,$ (1.1)
where the left-hand side represents the Schwarzian derivative. This derivative arises in various areas of mathematics, including classical complex analysis, one-dimensional dynamics, integrable systems, and conformal field theory [7].
Equation (1.1) admits the following six-dimensional Lie algebra
$\begin{array}[]{ccc}\displaystyle X_{1}=\frac{\partial}{\partial x},\qquad X_{2}=x\frac{\partial}{\partial x},\qquad X_{3}=x^{2}\frac{\partial}{\partial x},\\ \displaystyle X_{4}=\frac{\partial}{\partial y},\qquad X_{5}=y\frac{\partial}{\partial y},\qquad X_{6}=y^{2}\frac{\partial}{\partial y}.\end{array}$ (1.2)
In [8, 9], the adjoint equation method [4] was employed to construct the first integrals of (1.1), which will be discussed further. Through the integrals, one derives the general solution of (1.1) as
$y=\frac{1}{C_{1}x+C_{2}}+C_{3},$ (1.3)
where $C_{1}$, $C_{2}$, and $C_{3}$ are arbitrary constants.
The second equation under consideration is
$y^{\prime\prime}+2\,\frac{y^{\prime}+C{y^{\prime}}^{3/2}+{y^{\prime}}^{2}}{x-y}=0,$ (1.4)
where $C$ is an arbitrary constant. Notice that this constant cannot be eliminated by any equivalence transformation, i.e., without altering the group structure of the equation. This second-order ODE belongs to the Lie classification list [10, 11]. In his classification, Lie identified all second-order ODEs that admit nontrivial symmetries. Remarkably, most second-order ODEs whose solutions and integrals appear in standard handbooks are equivalent, up to point transformations, to equations from the Lie list.
According to the classification, equation (1.4) admits the tree-dimensional Lie algebra
$Y_{1}=\frac{\partial}{\partial x}+\frac{\partial}{\partial y},\qquad Y_{2}=x\frac{\partial}{\partial x}+y\frac{\partial}{\partial y},\qquad Y_{3}=x^{2}\frac{\partial}{\partial x}+y^{2}\frac{\partial}{\partial y}.$ (1.5)
Employing this symmetries and the Noether theorem, the general solution of (1.4) is derived [8] as
$y=\frac{1}{A(B-Ax)}+\frac{B-C}{A},$ (1.6)
where $A$ and $B$ are constants of integration. Equation (1.4) also has the singular solution
$y=ax+b,$ (1.7)
where $a$ and $b$ are constants, and $a$ must satisfy the condition
$a(a+C\sqrt{a}+1)=0.$ (1.8)
Notice that (1.6) is a particular case of the solution (1.3). Indeed, a connection between (1.1) and (1.4) can be established. [^1] For the case $C=0$, the connection between the solutions of these ODEs was noted long ago; see, e.g., [12, p. 15], which references the Mathematical Tripos, Part I Examination, 1911. By solving (1.4) for the constant $C$, one obtains
$\frac{(y-x)y^{\prime\prime}-2y^{\prime}\left(1+y^{\prime}\right)}{2{y^{\prime}}^{3/2}}=C.$ (1.9)
Differentiating (1.9) with respect to $x$ yields
$\frac{y-x}{2\sqrt{y^{\prime}}}\left(\frac{y^{\prime\prime\prime}}{y^{\prime}}-\frac{3}{2}\left(\frac{y^{\prime\prime}}{y^{\prime}}\right)^{2}\right)=0.$ (1.10)
This shows that (1.9) is a first integral of the Schwarz equation (1.1), with the multiplier
$\frac{y-x}{2\sqrt{y^{\prime}}}.$
Similar to the Lie classification [10] for ODEs, a group classification of second-order finite-difference schemes was carried out in [13]. As a result, symmetry-preserving finite-difference schemes were constructed for the equations from the Lie list, including schemes for equation (1.4). Such symmetry-preserving schemes are called invariant, and their construction theory has been well developed in recent decades [14, 15]. Among invariant schemes, there exist exact schemes, which are schemes whose solutions coincide exactly with the solutions of the corresponding differential equations at the nodes of the finite-difference mesh. Such schemes for ODEs were constructed, e.g., in [16, 17, 18]. An exact scheme for equation (1.4) was derived in [18], while for the Schwarz equation (1.1), it was obtained in [19, 20].
As discrete analogues for both equations (1.1) and (1.4) have been constructed, a natural question arises about the possibility of establishing a connection between them, similar to the one described above for the corresponding ODEs. In the following sections, we will demonstrate that there exists a finite-difference connection of a more complex nature, based on the application of Bäcklund-type finite-difference transformations.
The remainder of the paper is organized as follows. In Section 2, we provide a brief overview of the invariant finite-difference schemes for the Schwarz equation (1.1) and the second-order ODE (1.4). The connection between these schemes using Bäcklund-type finite-difference transformations is explored in Section 3. Finally, Section 4 discusses the implications of the established connection and presents conclusions based on the findings.
2 Invariant Finite-Difference Schemes for Eqs. (1.1) and (1.4)
We consider schemes on the following subset (a four-point stencil) of the finite-difference space
$\left(x_{n-1},x_{n},x_{n+1},x_{n+2},u_{n-1},u_{n},u_{n+1},u_{n+2}\right)$ (2.1)
Further, the symbol $x_{n}=x(n)$ (or $t_{n}=t(n)$) is used as an approximation for the independent variable, and $u_{n}=u(n)$ (or $y_{n}=y(n)$) is used as an approximation for the dependent variable at the point corresponding to the integer index $n$ on the finite-difference grid. The index $n$ is shifted left or right by the finite-difference shift operators ${\underset{-h}{S}}$ and ${\underset{+h}{S}}$, respectively. The finite-difference derivatives are defined in terms of shifts as
$\underset{+h}{D}=\frac{{\underset{+h}{S}}-1}{x_{n+1}-x_{n}},\qquad\underset{-h}{D}=\frac{1-{\underset{-h}{S}}}{x_{n}-x_{n-1}}.$
The operators ${\underset{+h}{S}}$ and ${\underset{-h}{S}}$ commute with the differentiations $\underset{\pm h}{D}$ in any order.
For brevity, we adopt Samarskii’s notation [21] and write the variables (2.1) as
$\left(x_{-},x,x_{+},x_{++},u_{-},u,u_{+},u_{++}\right).$ (2.2)
In general, for any function $f=f(n)$, we denote
$f_{-}={\underset{-h}{S}}(f),\quad f_{+}={\underset{+h}{S}}(f),\quad f_{++}={\underset{+h}{S}}^{2}(f).$
We also denote the finite-difference derivatives and difference mesh steps as
$u_{\bar{x}}=\underset{-h}{D}(u),\qquad u_{x}=\underset{+h}{D}(u),\qquad u_{x}^{+}=\underset{+h}{D}(u_{+})=\underset{+h}{S}(u_{x}),$
$h_{-}=x-x_{-},\qquad h_{+}=x_{+}-x,\qquad h_{++}=x_{++}-x_{+}.$
In contrast to the continuous case [1, 2], in variables (2.2), the prolongation formula for the generator
$X=\xi\frac{\partial}{\partial x}+\eta\frac{\partial}{\partial u}$ (2.3)
is defined through finite-difference shifts [14] as follows:
[…]
A finite-difference equation $F(x_{-},x,x_{+},x_{++},u_{-},u,u_{+},u_{++})=0$ is called invariant with respect to the symmetry generator (2.3) if
[…]
Here $[F]$ indicates that (2.4) is evaluated on solutions of $F=0$ and all its finite-difference consequences.
An invariant finite-difference scheme on the stencil (2.2) is typically written as a system of two equations [22, 14],
$\begin{array}[]{l}\Phi(x_{-},x,x_{+},x_{++},u_{-},u,u_{+},u_{++})=0,\\ \Omega(x_{-},x,x_{+},x_{++},u_{-},u,u_{+},u_{++})=0,\end{array}$ (2.5)
for which the invariance criteria holds:
[…]
for all symmetries $X_{i}$ admitted by the corresponding ODE. The first equation in (2.5) serves as a finite-difference approximation of the ODE, while the second equation, referred to as the mesh equation, vanishes in the continuous limit, reducing to the trivial identity $0\equiv 0$.
An invariant scheme for (1.4) was constructed by the authors in [18] and numerically investigated in [23]. It can be written in the form
$\theta\left(\frac{\sqrt{u_{\bar{x}}}}{\sqrt{(x-u_{-})(x_{-}-u)}}-\frac{x_{+}-u}{x_{-}-u}\frac{\sqrt{u_{x}}}{\sqrt{(x-u_{+})(x_{+}-u)}}\right)=\frac{C\sqrt{1+\varepsilon}\,(x-x_{-})u_{\bar{x}}}{(x-u_{-})(x_{-}-u)},$ (2.6)
$\frac{(x_{+}-x)^{2}u_{x}}{(x-u_{+})(x_{+}-u)}=\frac{(x-x_{-})^{2}u_{\bar{x}}}{(x-u_{-})(x_{-}-u)}=\varepsilon,$ (2.7)
where the small parameter $\varepsilon$ characterizes the density of the finite-difference mesh, and $\theta$ ($\theta\to 1$) is a parameter. In case
$\theta=\frac{\sqrt{1+\varepsilon}}{2}\left(|C|\sqrt{\varepsilon}+\sqrt{\varepsilon C^{2}+4}\right).$ (2.8)
the scheme becomes exact, [^2] To simplify the calculations, in the next section we assume that $\theta=\textrm{sgn}\,{C}$. As verified through previous calculations carried out by the authors, this assumption does not qualitatively affect the subsequent results, while it substantially simplifies the equations. i.e., its solutions at the grid nodes coincide with those of the corresponding ODE. As scheme (2.6), (2.7) is invariant, it admits the same algebra (1.5) as that of (1.4).
By means of the difference analogue of the Noether theorem [14], the following three first integrals [18] were derived for the scheme
$\begin{array}[]{c}\displaystyle J_{1}=\frac{u_{x}+1}{\sqrt{u_{x}(x-u)(x_{+}-u_{+})}}+\frac{C}{x_{+}-u}=A,\\ \displaystyle J_{2}=\frac{x_{+}u_{x}+u}{\sqrt{u_{x}(x-u)(x_{+}-u_{+})}}+\frac{Cx_{+}}{x_{+}-u}=B,\\ \displaystyle J_{3}=\frac{(x_{+}-x)(u_{+}-u)}{(x-u_{+})(x_{+}-u)}=\varepsilon.\end{array}$ (2.9)
Using these integrals, the scheme can be reduced to a Riccati difference equation, which is solved by standard methods.
For the Schwarz equation (1.1), an invariant finite-difference scheme, known as the Winternitz scheme, was proposed in [19, 20] and has the form
$\begin{array}[]{l}\displaystyle\frac{(y_{++}-y)(y_{+}-y_{-})}{(y_{++}-y_{+})(y-y_{-})}-K=0,\\ \displaystyle\frac{(t_{++}-t)(t_{+}-t_{-})}{(t_{++}-t_{+})(t-t_{-})}-K=0,\end{array}$ (2.10)
where $K$ is constant. This scheme admits the same six-dimensional algebra (1.2) as the Schwarz equation.
In [8] and [9] scheme (2.10) was integrated for arbitrary values of $K$, as well as schemes of a somewhat more general form. This was first done for a difference equation of odd order using the difference analogue of the adjoint equation method [4], which is applicable even when the equation does not possess a Lagrangian or a Hamiltonian.
A comparison of the general solutions of scheme (2.6), (2.7) and scheme (2.10) found in [18, 8, 9] leads to the following relation between the constants $K$ and $C$
$K=\frac{\varepsilon C^{2}+4}{\varepsilon+1}.$
In the case $K=4$ ($C^{2}=4$), scheme (2.10) is exact [8]. For simplicity, further we restrict ourselves to this case, as the calculations for it are the most concise. The cases $K<4$ and $K>4$ can be addressed in a similar manner.
For $K=4$, the adjoint equation method gives six difference integrals of scheme (2.10). Here, we present only the two integrals that we will need later, namely
$\begin{array}[]{r}\displaystyle\frac{4}{y_{+}-y_{-}}-\frac{1}{y_{+}-y}-\frac{1}{y-y_{-}}=\textrm{const},\\ \displaystyle\frac{4}{t_{+}-t_{-}}-\frac{1}{t_{+}-t}-\frac{1}{t-t_{-}}=\textrm{const}.\end{array}$ (2.11)
By means of the integrals, the general solution of scheme (2.10) is found:
$y_{n}=\frac{1}{c_{1}n+c_{2}}+c_{3},\qquad t_{n}=\frac{1}{c_{4}n+c_{5}}+c_{6},$ (2.12)
where $c_{1}$, $c_{2}$, …, $c_{6}$ are constants of integration.
The exact solution of scheme (2.6), (2.7) corresponding to the case $K=4$ is
$u_{n}=\frac{1}{A(B-Ax_{n})}+\frac{B-C}{A},\qquad x_{n}=\frac{\textrm{sgn}\,{C}\sqrt{1+\varepsilon}}{A\sqrt{\varepsilon}(\rho+n)}+\frac{B-\textrm{sgn}\,{C}}{A},$ (2.13)
where $\rho$ is constant.
Remark 2.1.
Solution (2.13) can be rewritten as
$u_{n}=\frac{\textsl{sgn}\,{C}\sqrt{\varepsilon}(\rho+n)}{A\left(\sqrt{\varepsilon}(\rho+n)-\sqrt{1+\varepsilon}\right)}+\frac{B-C}{A},\quad x_{n}=\frac{\textrm{sgn}\,{C}\sqrt{1+\varepsilon}}{A\sqrt{\varepsilon}(\rho+n)}+\frac{B-\textrm{sgn}\,{C}}{A}.$
Remark 2.2.
Using (2.13), one can reconstruct the remaining first integral of (2.6), (2.7) for $C^{2}=4$:
$J_{4}=\frac{1}{x_{+}-x-u_{+}+u}\left(u-x+{\textsl{sgn}\,{C}}\sqrt{u_{x}(u-x)(u_{+}-x_{+})}\right)-(n+1)=\rho.$
Apparently, this first integral cannot be obtained by means of the difference analogue of the Noether theorem due to its dependence on the index $n$, which in some sense expresses the ‘non-locality’ of the integral.
Remark 2.3.
As equation (1.4) has the singular solution (1.7), and scheme (2.6), (2.7) is exact, it should also possess a singular solution of a similar form. To verify this, substitute
$u_{n}=ax_{n}+b$ (2.14)
into (2.7), resulting in the equation
$x_{n+1}=\frac{1}{2a(\varepsilon+1)}\left\{\left((a-1)\sqrt{\varepsilon(a^{2}\varepsilon+2(\varepsilon+2)a+\varepsilon)}+\varepsilon+(\varepsilon a+2)a\right)x_{n}\right.\\ \left.+\,b\left(\sqrt{\varepsilon(a^{2}\varepsilon+2(\varepsilon+2)a+\varepsilon)}+\varepsilon(a-1)\right)\right\}.$
This is a linear equation solvable by standard methods. Substituting this into (2.6) yields
$a^{2}C\sqrt{2(\varepsilon+1)}\sqrt{(a-1)\sqrt{\varepsilon(\varepsilon a^{2}+2(\varepsilon+2)a+\varepsilon)}+\varepsilon+(\varepsilon a+2)a}\\ -\theta a(a+1)\left(\sqrt{\varepsilon(a^{2}\varepsilon+2(\varepsilon+2)a+\varepsilon)}-\varepsilon-(\varepsilon+2)a\right)=0.$ (2.15)
Taking into account (2.8), in the continuous limit $(\varepsilon\to 0)$, one gets
$a\,(a+C\sqrt{a}+1)+O(\sqrt{\varepsilon})=0,$
which corresponds to the condition (1.8). This can also be interpreted as a condition on the choice of $\varepsilon$: for the scheme to have the singular solution (2.14), the mesh density parameter $\varepsilon$ must satisfy (2.15).
3 Establishing a Finite-Difference Connection Through Bäcklund-type Transformations (Case $K=C^{2}=4$)
Suppose there exists a connection between scheme (2.6), (2.7) and scheme (2.10) through finite-difference differentiation, similar to the corresponding ODEs. To verify this, equation (2.6) is first rewritten in an integral form by solving for $C$:
$\frac{(x-u_{-})(x_{-}-u)}{(u-u_{-})\sqrt{1+\varepsilon}}\left(\frac{\sqrt{u_{\bar{x}}}}{\sqrt{(x-u_{-})(x_{-}-u)}}-\frac{x_{+}-u}{x_{-}-u}\frac{\sqrt{u_{x}}}{\sqrt{(x-u_{+})(x_{+}-u)}}\right)=C.$ (3.1)
Next, assuming that (3.1) can be reduced to a first integral of the Winternitz scheme (2.10), and given that (2.10) is symmetric with respect to $y$ and $t$, there should be another integral obtained from (3.1) by interchanging $u$ and $x$. As the calculations confirm, the following equation holds on the mesh (2.7)
$\frac{(u-x_{-})(u_{-}-x)}{(x-x_{-})\sqrt{1+\varepsilon}}\left(\frac{\sqrt{x_{\bar{u}}}}{\sqrt{(u-x_{-})(u_{-}-x)}}-\frac{u_{+}-x}{u_{-}-x}\frac{\sqrt{x_{u}}}{\sqrt{(u-x_{+})(u_{+}-x)}}\right)=\widetilde{C},$ (3.2)
where we have denoted
$x_{u}=\frac{x_{+}-x}{u_{+}-u},\qquad x_{\bar{u}}=\frac{x-x_{-}}{u-u_{-}},\qquad\widetilde{C}=\frac{C}{1+C\sqrt{\varepsilon(1+\varepsilon)}}.$
Thus, scheme (3.1), (3.2) has been obtained in integral form, and it is equivalent to scheme (2.6), (2.7).
Differentiating the scheme by applying the operator $\underset{+h}{D}$ to (3.1) and (3.2), one derives the four-point scheme
$\begin{array}[]{c}\displaystyle\frac{(x_{+}-u)(x-u_{+})}{u_{+}-u}\left(\frac{\sqrt{u_{x}}}{\sqrt{(x_{+}-u)(x-u_{+})}}-\frac{x_{++}-u_{+}}{x-u_{+}}\frac{\sqrt{u_{x}^{+}}}{\sqrt{(x_{+}-u_{++})(x_{++}-u_{+})}}\right){}\\ \displaystyle{}-\frac{(x-u_{-})(x_{-}-u)}{u-u_{-}}\left(\frac{\sqrt{u_{\bar{x}}}}{\sqrt{(x-u_{-})(x_{-}-u)}}-\frac{x_{+}-u}{x_{-}-u}\frac{\sqrt{u_{x}}}{\sqrt{(x-u_{+})(x_{+}-u)}}\right)=0,\\ {}\hfil\\ \displaystyle\frac{(u_{+}-x)(u-x_{+})}{x_{+}-x}\left(\frac{\sqrt{x_{u}}}{\sqrt{(u_{+}-x)(u-x_{+})}}-\frac{u_{++}-x_{+}}{u-x_{+}}\frac{\sqrt{x_{u}^{+}}}{\sqrt{(u_{+}-x_{++})(u_{++}-x_{+})}}\right){}\\ \displaystyle{}-\frac{(u-x_{-})(u_{-}-x)}{x-x_{-}}\left(\frac{\sqrt{x_{\bar{u}}}}{\sqrt{(u-x_{-})(u_{-}-x)}}-\frac{u_{+}-x}{u_{-}-x}\frac{\sqrt{x_{u}}}{\sqrt{(u-x_{+})(u_{+}-x)}}\right)=0.\end{array}$ (3.3)
The first equation of (3.3) approximates the Schwarz equation, while the second one serves as a mesh equation. Clearly, (3.1) and (3.2) are first integrals of (3.3), as system (3.3) was obtained by differentiating them.
It might be expected that scheme (3.3) could be transformed into the Winternitz scheme (2.10) via some point transformation; however, this is not the case. It can be verified that scheme (3.3) admits only the three-dimensional subalgebra (1.5) of the six-dimensional algebra (1.2), which is admitted by the Winternitz scheme. This implies that scheme (3.3) cannot be transformed into scheme (2.10) by a point transformation, and therefore, more general transformations must be considered. Note also that scheme (3.3) remains exact on the subset (2.13) of the solutions of the Winternitz scheme.
To proceed, consider Bäcklund-type transformations, starting with the differential case. Recall that the equation
$\mathcal{B}\left(x,u,y,u^{\prime},y^{\prime},u^{\prime\prime},y^{\prime\prime}\right)=0$ (3.4)
defines a Bäcklund transformation that relates equations
$\frac{u^{\prime\prime\prime}}{u^{\prime}}-\frac{3}{2}\left(\frac{u^{\prime\prime}}{u^{\prime}}\right)^{2}=0$ (3.5)
and
$\frac{y^{\prime\prime\prime}}{y^{\prime}}-\frac{3}{2}\left(\frac{y^{\prime\prime}}{y^{\prime}}\right)^{2}=0$ (3.6)
if the compatibility of (3.4) and (3.5) implies (3.6), and the compatibility of (3.4) and (3.6) implies (3.5). In this particular case, both (3.5) and (3.6) are Schwarz equations, and a transformation of this type is referred to as an auto-Bäcklund transformation. By analyzing the compatibility conditions of (3.4), (3.5), and (3.6), it can be demonstrated that an equation of the form
$\mathcal{B}\left(\frac{u^{\prime\prime}}{{u^{\prime}}^{3/2}},u-\frac{2{u^{\prime}}^{2}}{u^{\prime\prime}},\frac{y^{\prime\prime}}{{y^{\prime}}^{3/2}},y-\frac{2{y^{\prime}}^{2}}{y^{\prime\prime}}\right)=0,$ (3.7)
where the arguments of $\mathcal{B}$ are first integrals of (1.1), defines an auto-Bäcklund transformation. As a straightforward example of such a transformation, one can consider
$\mathcal{B}=u-\frac{2{u^{\prime}}^{2}}{u^{\prime\prime}}+\alpha\left(\frac{y^{\prime\prime}}{{y^{\prime}}^{3/2}}\right)=0,$ (3.8)
where $\alpha$ is some nonzero constant which can be expressed in terms of integration constants of (3.5) and (3.6).
Assuming that scheme (3.3) corresponds to equation (3.5) and scheme (2.10) corresponds to equation (3.6), and using the known first integrals of these schemes, one can construct a difference analogue of the transformation (3.7). It is important to note that in the difference case, two equations, $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$, are needed to connect independent first integrals from both schemes, as the systems (3.3) and (2.10) include mesh equations. An example of such a transformation is
$\begin{array}[]{c}\displaystyle\mathcal{B}_{1}=\frac{(x-u_{-})(x_{-}-u)}{u-u_{-}}\left(\frac{\sqrt{u_{\bar{x}}}}{\sqrt{(x-u_{-})(x_{-}-u)}}-\frac{x_{+}-u}{x_{-}-u}\frac{\sqrt{u_{x}}}{\sqrt{(x-u_{+})(x_{+}-u)}}\right)\\ \displaystyle{}+\alpha_{1}\left(\frac{4}{y_{+}-y_{-}}-\frac{1}{y_{+}-y}-\frac{1}{y-y_{-}}\right)=0,\\ {}\displaystyle\mathcal{B}_{2}=\frac{(u-x_{-})(u_{-}-x)}{x-x_{-}}\left(\frac{\sqrt{x_{\bar{u}}}}{\sqrt{(u-x_{-})(u_{-}-x)}}-\frac{u_{+}-x}{u_{-}-x}\frac{\sqrt{x_{u}}}{\sqrt{(u-x_{+})(u_{+}-x)}}\right)\\ \displaystyle{}+\alpha_{2}\left(\frac{4}{t_{+}-t_{-}}-\frac{1}{t_{+}-t}-\frac{1}{t-t_{-}}\right)=0,\end{array}$ (3.9)
where $\alpha_{1}$ and $\alpha_{2}$ are nonzero constants which can be expressed in terms of the integration constants of the schemes. Here, four integrals (2.11), (3.1), and (3.2) are employed, though this transformation is not the only possible one. It is easy to verify that system (3.3) follows from the compatibility conditions of (2.10) and (3.9), and system (2.10) follows from the compatibility conditions of (3.3) and (3.9). In other words, a Bäcklund-type difference transformation relating schemes (2.10) and (3.3) was constructed.
Based on the above construction, one formulates the following more general
Proposition.
Given two schemes for ODEs of the same order $k>1$ with at least two known first integrals each, a finite-difference Bäcklund-type transformation can be constructed utilizing these integrals to relate the schemes. If both schemes approximate the same ODE, this transformation reduces to an auto-Bäcklund transformation in the continuous limit.
It is worth noting that discrete Bäcklund-type transformations have previously been applied to differential-difference equations [15] and to problems related to the linearization of finite-difference schemes [24].
Remark 3.1.
As the general solutions (2.12) and (2.13) of schemes (3.3) and (2.10) are known, substituting these solutions,
$x_{\pm}=x_{\pm}(n),\quad t_{\pm}=t_{\pm}(n),\quad u_{\pm}=u_{\pm}(n),\quad y_{\pm}=y_{\pm}(n),\quad\cdots,$ (3.10)
allows explicit relations between $x$, $u$, $t$, and $y$ to be expressed in terms of the index $n$ and integration constants. The expressions for transformation (3.9) are quite cumbersome and are not presented here.
4 Conclusions
An example was considered in which the second-order ODE (1.4) is transformed into a form where it becomes a first integral (1.9) of the Schwarz equation (1.1). It was shown that, in the discrete case, this structure does not hold for the known finite-difference schemes.
By differentiating the exact scheme for the second-order equation, written in terms of the first integrals (3.1) and (3.2), a scheme for equation (1.1) is obtained that does not coincide with the known exact scheme (2.10). Instead, the scheme (3.3) is derived, which is exact only for a subset of solutions of (3.1), (3.2) and admits only the subalgebra (1.5) of the six-dimensional algebra (1.2). This implies that the two schemes for the Schwarz equation cannot be related by a point transformation.
Nevertheless, a connection between schemes (3.3) and (2.10) can be established through a finite-difference transformation (3.9) of the Bäcklund type, utilizing the known first integrals of these two schemes. In the continuous limit, both schemes (2.10) and (3.3) reduce to the Schwarz equation (up to some multipliers), while the discrete transformation (3.9) simplifies to an auto-Bäcklund transformation, mapping the Schwarz equation onto itself.
The results suggest that multiple exact schemes for the same ODE may exist, connected by point transformations (for isomorphic Lie algebras) or more complex transformations like discrete Bäcklund-type transformations.
Acknowledgements
This research was partially supported by Suranaree University of Technology and the NSRF via the Program Management Unit for Human Resources & Institutional Development, Research and Innovation (PMU-B) (Grant B13F660067). E.I.K. is also grateful to Prof. S. V. Meleshko for support and fruitful discussions during the current postdoctoral fellowship (Full-time66/06/2023).